Combined, there are 188 ​Asians, Africans,​ Europeans, and Americans in a village. The number of Asians exceeds the number of Africans and Europeans by 72. The difference between the number of Europeans and Americans is 13. If the number of Africans is​ doubled, their population exceeds the number of Europeans and Americans by 27. Determine the number of​ Asians, Africans,​ Europeans, and Americans in this village.
Answer by ikleyn(52788)

.
Combined, there are 188 ​Asians, Africans,​ Europeans, and Americans in a village.
The number of Asians exceeds the number of Africans and Europeans by 72.
The difference between the number of Europeans and Americans is 13.
If the number of Africans is​ doubled, their population exceeds the number of Europeans and Americans by 27.
Determine the number of​ Asians, Africans,​ Europeans, and Americans in this village.
~~~~~


x = As

y = Af

z = Eu

t = Am


The system of equation


    x + y + z + t = 188

    x - y - z     =  72

            z - t =  13

       2y - z - t =  27


Next, I use the free of charge online solver  www.reshish.com

to get the solution to this system.


ANSWER.  x= 125 (Asians);  y= 30 (Africans);  z= 23 (Europeans);  t= 10 (Americans).

Solved.